# DreamBench++: A Human-Aligned Benchmark for Personalized Image Generation

### v1

@@ 3.1 Experimental Setup
Human Annotators  We employ 7 human annotators to score for each instance in DreamBench++ to obtain ground truth human preference data. We provide human annotators with sufficient training to ensure they fully understand the personalized image generation task and can provide unbiased and discriminating scores. The scoring task and scheme given to humans are identical to those used for GPT, as described in Section 2. The GPT results and human results are isolated to avoid hindsight bias. Additionally, we ensure that each instance is rated by at least two humans to reduce noise.
Figure 7: A qualitative study of different methods on DreamBench++. We demonstrate the generation quality of different methods on our DreamBench++, including animals, humans, objects, and style, with photo and non-photo-realistic examples. The blue block highlights fine-tuning-based methods, and the green block highlights encoder-based methods. Instances above the dotted line are evaluated for subject preserving, and instances below are evaluated for style preserving.
@@ 3.2 Main Results
Quantitative Analysis  Table 1 shows the overall evaluation results, including human and GPT-4o rating scores. From the results, we observe that: i) DreamBench++ aligns better with humans than DINO or CLIP models. Driven by our dedicatedly-designed prompts, GPT-4o used by DreamBench++ yields impressive alignment with humans. This is because humans and DreamBench++ are all advanced in evaluating facial and textural characters and producing scores with a balanced consideration of all aspects. ii) DINO-I and CLIP-I yield significant divergence from humans in evaluating concept preservation. This could be because DINO/CLIP scores show a preference for images that preserve shapes or overall styles, as shown in Fig. 6. For example, methods proficient in posture and detailed texture transformation, such as DreamBooth-L, rank lower than humans and DreamBench++. iii) Traditional CLIP-T scores are as effective as DreamBench++ in evaluating prompt following, showing strong alignment with humans.
@@
Animal Human Object Style Overall Photorealistic Style Imaginative Overall
@@
Qualitative Analysis  Fig. 7 shows selected qualitative results on DreamBench++, which helps provide an intuitive understanding of evaluated models. With a more comprehensive and diverse collection of images, we have discovered numerous intriguing characteristics of these generation methods that were not apparent on existing datasets such as DreamBench. Specifically, we observe that: i) Fine-tuning-based methods outperform encoder-based methods on images containing more subject-oriented information, such as an animal or object, as they preserve more intricate details in the generated images. However, for images containing a person, fine-tuning-based methods often fail to preserve facial and clothing features. This suggests that the personalized generation of human images is more demanding for visual concept preservation than for textual following capabilities, which is an advantage for encoder-based methods. ii) However, for style-oriented cases when subject details are less critical, encoder-based methods perform better than fine-tuning-based methods. This further highlights the strengths of encoder-based methods in that they are more adept at recognizing and extracting high-level visual semantics, including overall shape, style, and thematic features.
Leaderboard  Table 2 shows the detailed results of the leaderboard with respect to the concept and prompt categories defined in Section 2. Note that: i) the human category shows the lowest average score of 0.482, which is -0.204 lower than the highest average score of animal. This category is very challenging in terms of concept preservation because humans are very sensitive to facial details, and many works are conducted specifically on it [105, 111, 100, 114]. ii) The object is also a relatively difficult category because the characteristics of each same or different object are often very diverse. In contrast, animals within the same category often share a large degree of visual similarity. iii) There exists a negative correlation between concept preservation and prompt following. The primary aim of personalized T2I evolution is to identify the Pareto optimum that balances both factors.
@@
Method TI [28] DreamBooth [86] DreamBooth-L [86] BLIP-D [55] Emu2 [95] IP-Adapt.-P [115] IP-Adapt. [115]
@@
3.3 Ablation Study
Table 3 shows the ablation study of the prompt design influences on alignment with humans. From the results, we observe that: i) The proposed prompt designs are all necessarily effective, demonstrating the superiority of the prompting method in DreamBench++. For example, removing the proposed internal thinking leads to a significant drop, indicating the effectiveness of self-alignment for alignment with humans. ii) The capability of the multimodal GPT used is scalable. This shows that DreamBench++ has the potential to be improved with stronger GPT models in the future. iii) Some human prior knowledge, such as reminding the GPT not to consider background when assessing visual concept preservation, leads to performance degradation.
Table 4: The human alignment degree among different evaluation metrics, measured by Krippendorff’s alpha value. H, G, D, and C represent Human, GPT-4o, DINO Score, and CLIP Score, respectively. H-H value is also calculated to illustrate human self-alignment.
Method T2I Model Concept Preservation $\bm{\mathrm{Kd}_{\bar{O}}}$ Prompt Following $\bm{\mathrm{Kd}_{\bar{O}}}$
@@ H-H G-H D-H C-H H-H G-H C-H
$\bm{\mathrm{Ratio}_{\bar{O}}}$ 100% 79.64% 25.54% 3.34% 100% 93.18% 42.48%
4 Discussions
4.1 Is DreamBench++ aligned with humans?
Table 4 shows a more rigorous study of human alignment level using the mean Krippendorff’s alpha value [34]. The results show that DreamBench++ is a highly human-aligned benchmark. Notably, DreamBench++ achieves 79.64% and 93.18% evaluation consistency with human’s evaluation in concept preservation and prompt following capabilities, respectively. This result is +54.1% and +50.7% higher than traditional DINO and CLIP metrics.
Table 5: DreamBench [86] and DreamBench++ results comparison with traditional metrics. Unlike DreamBench, DreamBench++ uses one single reference image per instance; thus, the training steps and learning rate of $\bullet$ fine-tuning-based methods are slightly reduced to avoid overfitting.
Method T2I Model DreamBench [86] DreamBench++
@@ 4.1 Is DreamBench++ aligned with humans?
4.2 Is data diversity necessary?
To investigate the necessity of diverse and rich evaluation data, we compare the results on DreamBench and DreamBench++ using DINO and CLIP metrics. Table 5 demonstrates that diverse data in DreamBench++ is necessary to avoid biased evaluation. Although the results are overall consistent, especially for encoder-based methods, fine-tuning-based methods such as TI and DreamBooth both show a notable score drop. We argue that this is due to the reduced training steps and learning rate limited by the single-image fine-tuning, which may make the model not fully converged. However, unlike other encoder-based methods, Emu2 shows a significant score drop. Emu2 performs well only on natural images and simple text, while it cannot generate valid results when texts are complex or stylized or with anime reference images, as shown in Fig. 8.
Figure 8: Case study of successful and failure case on DreamBench++ with Emu2.
@@ 4.2 Is data diversity necessary?
Method TI [28] DreamBooth [86] DreamBooth-L [86] BLIP-D [55] Emu2 [95] IP-Adapt.-P [115] IP-Adapt. [115]
@@
4.3 Can we use free lunch to improve DreamBench++ evaluation?
Table 6 shows the result of utilizing free lunch techniques, including chain-of-thought (CoT) [108] and In-Context Learning (ICL) [4, 12]. CoT indicates that GPT-4 articulates its reasoning process before scoring, and ICL indicates GPT-4o is provided with human-written few-shot examples.
@@ 5 Related Works
Personalized Image Generation aims to preserve concept consistency while accommodating the diverse contexts suggested by the instructions. In general, it can be traced back to early efforts on pixel-to-pixel (Pix2Pix) translation where the personalization orientation is free-form texts [11, 99, 75] or predefined translation across styles, seasons, species, or plants, etc [46, 127, 118, 106, 87]. Modern efforts go beyond Pix2Pix translation toward a free-form image generation conditioned on both reference images and prompts. Some works focus on fine-tuning techniques that turn a general T2I model into a specialist personalization model [28, 86, 53, 92, 74] using LoRA [42] or contrastive learning [117, 35], learning the subject or style information by reconstructive autoencoding [102, 23]. However, the necessity to fine-tune for new subjects limits their scalability. In contrast, encoder-based methods can generate subject-guided or style-guided images or edit images following prompts with one shot. Encoder- or adapter-based methods [125, 115, 109, 55, 48, 29, 16, 103, 105] train an encoder to encode the conditional image into embeddings, which are integrated into cross-attention mechanism in the diffusion process [40, 93, 68, 94]. Adapter-free methods [66, 64, 37, 36, 11] extract the information, such as attention maps [36, RBModulation24] from reference images, and fuse them into the image generation process. Furthermore, multimodal large language models (MLLMs) that are trained on extensive multimodal sequences can also serve as general foundation models [24, 95, 73, 30].
@@
Benchmarking Image Generation involves a variety of metrics that focus on different aspects. Inception Score [89] and FID [39] judge image quality, while LPIPS [120], DreamSim [26], CLIP-I [81], and DINO Score [13] measure perceptual similarity. In text-guided generation, prompt-image alignment can be assessed by CLIP-T [81], CLIPScore [38], and BLIP Score [56, 57]. However, these metrics often fall short of reflecting human perception. To address this, human-aligned metrics [52, 112, 54] have been introduced, offering a more perceptive evaluation. Yet, they face limitations in scaling with the pace of new model developments. Thus, the necessity for automated and sustainable evaluation methods has emerged, with some [112, 60, 32] leveraging reward-model-based methods to encode human preferences, while others [51, 18, 110, 121, 43, 65] use multimodal  [12, 84, 5, 62] to automate the process and better mirror human tastes. While MLLM-based methods show promise in aligning with human preferences [121, 110, 45, 19], automated personalized evaluation remains an unresolved issue. VIEScore [51] assesses image generation quality by prompting GPT-4V [70] and LLaVA [62, 61], but is limited to four models in subject-driven tasks and obtains suboptimal results. Meanwhile, Dreambench [86], a common benchmark for personalized generative evaluation, only consists of 30 simple objects and lacks diversity comprehensiveness.
@@ 6 Conclusions
This paper introduces DreamBench++, a human-aligned personalized image generation benchmark. Extensive and comprehensive experiments have demonstrated our advantages in dataset diversity and complexity, as well as the alignment of our evaluation metrics with human preferences. In addition, we offer insights into prompt design for advanced multimodal GPTs, emphasizing the potential and challenges of enhancing GPT evaluation through chain-of-thought prompting and in-context learning. Our work aims to support future research on personalized image generation by providing a human-aligned benchmark and heuristics in utilizing advanced multimodal GPTs in visual evaluation.
Appendix A Implementation Details
@@
Textual Inversion [28] SD v1.5 4 5e-4 3000 1 5e-4 3000
Dreambooth [86] SD v1.5 1 2.5e-6 1000 1 2.5e-6 250
Dreambooth LoRA [86] SDXL v1.0 4 5e-5 500 1 5e-5 500
@@
BLIP-Diffusion [55] SD v1.5 over-exposure, under-exposure, saturated, duplicate, out of frame, lowres, cropped,
@@
IP-Adapter ViT-G [115] SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
IP-Adapter-Plus ViT-H [115] SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
We dedicate significant effort to tuning hyperparameters to ensure that the performance of each method on DreamBench is consistent with results reported in original papers. As shown in Table 9, the results of our reproduction are comparable to or even better than the official results reported by the original papers.
@@
Appendix B Limitation & Future Work
Human-aligned evaluation & benchmarking is an emerging but challenging direction, and we have only made preliminary attempts at personalized image generation. Moreover, our evaluation results heavily rely on the advancements of multimodal large language models and require carefully designed system prompts. We believe that as visual world models continue to develop, the evaluation performance will be further optimized. Our future work will focus on more applications with human-aligned evaluation, such as 3D generation [78, 79, 63, 27], video generation [41, 91, 9], autonomous driving [33, 58, 119], and even embodied visual intelligence [59, 10, 25, 80, 123, 31].

### v2

@@ 3.1 Experimental Setup
Human Annotators  We employ 7 human annotators to score each instance in DreamBench++ to obtain ground truth human preference data. We provide human annotators with sufficient training to ensure they fully understand the personalized T2I generation task and can provide unbiased and discriminating scores. The scoring task and scheme given to humans are identical to those used for GPT, as described in Section 2. The GPT results and human results are isolated to avoid hindsight bias. Additionally, we ensure that each instance is rated by at least two humans to reduce noise.
@@ 3.2 Main Results
Quantitative & Qualitative Analysis  Table 1 shows the overall evaluation results, including human and GPT-4o rating scores. The results show that: i) DreamBench++ aligns better with humans than DINO or CLIP models. Driven by our dedicatedly-designed prompts, GPT-4o used by DreamBench++ yields impressive alignment with humans. This is because humans and DreamBench++ are all advanced in evaluating facial and textural characters and producing scores with a balanced consideration. ii) DINO-I and CLIP-I yield significant divergence from humans in evaluating concept preservation. This could be because DINO/CLIP scores show a preference for images that preserve shapes or overall styles (see Fig. 6). iii) Traditional CLIP-T scores are as effective as DreamBench++ in evaluating prompt following, showing strong alignment with humans. See qualitative results in Appendix A for an intuitive understanding of evaluated models.
@@
Animal Human Object Style Overall Realistic Style Imaginative Overall
@@
Leaderboard  Table 2 shows the leaderboard results with respect to the concept and prompt categories defined in Section 2. Note that: i) the human category shows the lowest average score of 0.482, which is -0.204 lower than the highest average score of animal. This category is very challenging in terms of concept preservation because due to facial details, and many works are conducted specifically on it (Wang et al., 2024b; Xiao et al., 2023; Valevski et al., 2023; Yan et al., 2023). ii) The object is also a relatively difficult category due to object diversity. In contrast, animals within the same category often share a strong visual similarity. iii) There exists a negative correlation between concept preservation and prompt following. The primary aim of personalized T2I evolution is to identify the Pareto optimum that balances both. iv) The CP/PF ratio reflects the over-fitting issue. A higher ratio indicates excessive adherence to the reference at the cost of prompt alignment, while a lower ratio suggests better prompt adherence but weaker concept preservation.
@@
Method TI DreamBooth DreamBooth-L BLIP-D Emu2 IP-Adapt.-P IP-Adapt.
@@
Abaltion Study  Table 3 shows the ablation study of the prompt design influences on alignment measured by mean Krippendorff’s alpha value (Hayes & Krippendorff, 2007). We observe that: i) The proposed prompt designs are all necessarily effective, demonstrating the superiority of the prompting method in DreamBench++. For example, removing the proposed internal thinking leads to a significant drop, indicating the effectiveness of self-alignment. ii) The capability of the multimodal GPT used is scalable. This shows that DreamBench++ has the potential to be improved in the future. iii) Some human prior knowledge, such as reminding the GPT not to consider background when assessing visual concept preservation, leads to performance degradation.
4 Discussions
4.1 Is DreamBench++ aligned with humans?
Table 4 shows a more rigorous study of human alignment level using the mean Pearson correlation value (Pearson, 1895). The results show that DreamBench++ is a highly human-aligned benchmark. Notably, DreamBench++ achieves 83.31% and 98.71% evaluation consistency with human’s evaluation in concept preservation and prompt following capabilities, respectively. This result is +32.59% and +37.23% higher than traditional DINO and CLIP metrics.
Table 4: The human alignment degree among different evaluation metrics, measured by Pearson correlation value. H, G, D, and C represent Human, GPT-4o, DINO Score, and CLIP Score, respectively. H-H value is also calculated to illustrate human self-alignment.
Method T2I Model Concept Preservation $\bm{\mathrm{PC}_{\bar{O}}}$ Prompt Following $\bm{\mathrm{PC}_{\bar{O}}}$
@@ H-H G-H D-H C-H H-H G-H C-H
$\bm{\mathrm{Ratio}_{\bar{O}}}$ 100% 83.31% 50.72% 60.98% 100% 98.17% 61.48%
4.2 Is data diversity necessary?
To assess the importance of diverse data, we compare results on DreamBench and DreamBench++ using DINO and CLIP metrics. Table 5 shows that the diverse data in DreamBench++ is key to unbiased evaluation. While overall results are consistent, TI, DreamBooth, and Emu2 show notable score drops. These methods perform well on natural images and simple text but struggle with complex or stylized prompts and anime references, see Fig. 7.
Table 5: DreamBench and DreamBench++ results comparison with traditional metrics. Unlike DreamBench, DreamBench++ uses a single reference image per instance; thus, the training steps and learning rate of $\bullet$ fine-tuning-based methods are slightly reduced to avoid overfitting.
Method T2I Model DreamBench DreamBench++
@@ 4.2 Is data diversity necessary?
Figure 7: Case study of successful and failure case on DreamBench++. The left images are reference images and the right images are results generated by Emu2, DreamBooth, and TI.
4.3 Can we use free lunch to improve DreamBench++ evaluation?
Table 6 shows the result of utilizing free lunch techniques, including chain-of-thought (CoT) (Wei et al., 2022) with GPT-4 articulating reasoning process and In-Context Learning (ICL) (Alayrac et al., 2022; Brown et al., 2020) with human-written few-shot examples.
@@ 4.3 Can we use free lunch to improve DreamBench++ evaluation?
Method TI DreamBooth DreamBooth-L BLIP-D Emu2 IP-Adapt.-P IP-Adapt.
@@
4.4 How transferable is DreamBench++ across MLLMs?
Table 7 reveals: i) DreamBench++ demonstrates robust generalization across different multimodal LLMs. The consistency ratios remain stable across various models, with Claude-3.5 achieving 74.00% and GPT-4o variants ranging from 67.22% to 84.23%. ii) The performance of MLLMs is continuously improving, as shown by the increasing Human-MLLM consistency in newer versions of GPT-4o (from 79.47% to 84.23%). This suggests that the accuracy of DreamBench++ evaluations may further improve as models advance.
Table 7: Comparison of Human-MLLM consistency ratios. Ratio: the overall alignment between model evaluations and human evaluations, calculated as the average proportion of H-H consistency.
Method TI DreamBooth DreamBooth-L BLIP-D Emu2 IP-Adapt.-P IP-Adapt. Ratio
T2I Model SD v1.5 SD v1.5 SDXL v1.0 SD v1.5 SDXL v1.0 SDXL v1.0 SDXL v1.0
H-H (Ground Truth) 0.685 0.647 0.656 0.613 0.746 0.602 0.591 100%
GPT-4o-2024-05-13 (DreamBench++) 0.544 0.596 0.641 0.362 0.669 0.366 0.458 79.47%
GPT-4o-2024-08-06 0.558 0.625 0.645 0.383 0.708 0.427 0.502 84.23%
GPT-4o-mini-2024-07-18 0.496 0.575 0.538 0.274 0.702 0.238 0.289 67.22%
Claude-3.5-Sonnet-20241022 0.500 0.583 0.623 0.354 0.625 0.278 0.427 74.00%
Gemini-1.5-Pro-001 0.487 0.547 0.514 0.267 0.653 0.253 0.202 63.04%
@@ 5 Related Works
aims to preserve concept consistency while accommodating the diverse contexts suggested by the instructions. In general, it can be traced back to early efforts on pixel-to-pixel (Pix2Pix) translation where the personalization orientation is free-form texts (Brooks et al., 2023; Tumanyan et al., 2023; Parmar et al., 2023) or predefined translation across styles, seasons, species, or plants, etc (Isola et al., 2017; Zhu et al., 2017; Zhang et al., 2023a; Wang et al., 2018; Saharia et al., 2022a). Modern efforts go beyond Pix2Pix translation toward a free-form image generation conditioned on both reference images and prompts. Some works focus on fine-tuning techniques that turn a general T2I model into a specialist personalization model (Gal et al., 2023a; Ruiz et al., 2023; Kumari et al., 2023; Sohn et al., 2023; Park et al., 2024) using LoRA (Hu et al., 2022) or contrastive learning (Zhang et al., 2022; He et al., 2020), learning the subject or style information by reconstructive autoencoding (Vincent et al., 2008; Dong et al., 2023). However, the necessity to fine-tune for new subjects limits their scalability. In contrast, encoder-based methods can generate subject-guided or style-guided images or edit images following prompts with one shot. Encoder- or adapter-based methods (Zheng et al., 2024; Ye et al., 2023; Wei et al., 2023; Li et al., 2023a; Jia et al., 2023; Gal et al., 2023b; Chen et al., 2023b; Wang et al., 2024a; b) train an encoder to encode the conditional image into embeddings, which are integrated into cross-attention mechanism in the diffusion process (Ho et al., 2020; Song et al., 2021a; Nichol & Dhariwal, 2021; Song et al., 2021b). Adapter-free methods (Lv et al., 2024; Liu et al., 2023d; Hertz et al., 2023b; a; Brooks et al., 2023) extract the information, such as attention maps (Hertz et al., 2023a) from reference images, and fuse them into the image generation process. Furthermore, multimodal large language models (MLLMs) that are trained on extensive multimodal sequences can also serve as general foundation models (Dong et al., 2024; Sun et al., 2024a; Pan et al., 2024; Ge et al., 2024). Additionally, some works (Wang et al., 2024b; Valevski et al., 2023; Yan et al., 2023; Ye et al., 2023; Xiao et al., 2023) focus on facial feature preservation.
Benchmarking Image Generation
@@
involves a variety of metrics that focus on different aspects. Inception Score (Salimans et al., 2016) and FID (Heusel et al., 2017) judge image quality, while LPIPS (Zhang et al., 2018), DreamSim (Fu et al., 2023), CLIP-I (Radford et al., 2021), and DINO Score (Caron et al., 2021) measure perceptual similarity. In text-guided generation, prompt-image alignment can be assessed by CLIP-T (Radford et al., 2021), CLIPScore (Hessel et al., 2021), and BLIP Score (Li et al., 2022; 2023b). However, these metrics often fall short of reflecting human perception. To address this, human-aligned metrics (Ku et al., 2024; Xu et al., 2023; Lee et al., 2023) have been introduced, offering a more perceptive evaluation. Yet, they face limitations in scaling with the pace of new model developments. Thus, the necessity for automated and sustainable evaluation methods has emerged, with some (Xu et al., 2023; Liang et al., 2023b; Guo et al., 2024) leveraging reward-model-based methods to encode human preferences, while others (Ku et al., 2023; Cho et al., 2023; Wu et al., 2024b; Zhang et al., 2023c; Hu et al., 2023; Lu et al., 2023) use multimodal  (Brown et al., 2020; Reid et al., 2024; Anil et al., 2023; Liu et al., 2023b) to automate the process and better mirror human tastes. While MLLM-based methods show promise in aligning with human preferences (Zhang et al., 2023c; Wu et al., 2024b; Huang et al., 2023; Cho et al., 2024), automated personalized evaluation remains an unresolved issue. VIEScore (Ku et al., 2023) assesses image generation quality by prompting GPT-4V (OpenAI, 2023) and LLaVA (Liu et al., 2023b; a), but is limited to four models in subject-driven tasks and obtains suboptimal results. Meanwhile, Dreambench (Ruiz et al., 2023), a common benchmark for personalized generative evaluation, only consists of 30 simple objects and lacks diversity comprehensiveness.
@@ 6 Conclusions
This paper introduces DreamBench++, a human-aligned personalized image generation benchmark. Extensive and comprehensive experiments show significant advantages in dataset diversity and complexity, along with metrics that align with human preferences. In addition, we offer insights into prompt design for advanced multimodal GPTs, emphasizing the potential and challenges of enhancing GPT evaluation through chain-of-thought prompting and in-context learning. Our work aims to support future research on personalized image generation by providing a human-aligned benchmark and heuristics in utilizing advanced multimodal GPTs in visual evaluation.
Appendix A Qualitative Analysis
With a more comprehensive and diverse collection of images, we have discovered numerous intriguing characteristics of these generation methods, as illustrated in Fig. 9, that were not apparent on existing datasets such as DreamBench. Specifically, we observe that: i) Fine-tuning-based methods outperform encoder-based methods on images containing more subject-oriented information, such as an animal or object, as they preserve more intricate details in the generated images. However, for images containing a person, fine-tuning-based methods often fail to preserve facial and clothing features. This suggests that the personalized generation of human images is more demanding for visual concept preservation than for textual following capabilities, which is an advantage for encoder-based methods. ii) However, for style-oriented cases when subject details are less critical, encoder-based methods perform better than fine-tuning-based methods. This further highlights the strengths of encoder-based methods in that they are more adept at recognizing and extracting high-level visual semantics, including overall shape, style, and thematic features.
Appendix B Implementation Details
@@
Textual Inversion SD v1.5 4 5e-4 3000 1 5e-4 3000
Dreambooth SD v1.5 1 2.5e-6 1000 1 2.5e-6 250
Dreambooth LoRA SDXL v1.0 4 5e-5 500 1 5e-5 500
@@
BLIP-Diffusion SD v1.5 over-exposure, under-exposure, saturated, duplicate, out of frame, lowres, cropped,
@@
IP-Adapter ViT-G SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
IP-Adapter-Plus ViT-H SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
We dedicate significant effort to tuning hyper-parameters to ensure that the performance of each method on DreamBench is consistent with results reported in original papers. Table 10 shows the results of our reproduction are comparable to or even better than the official results.
@@
Appendix C Additional Discussions
C.1 Are multiple images for each instance necessary?
In practice, multiple reference images are unnecessary for personalized image generation: i) The limited availability of reference images during daily usage makes single-image personalization more relevant. ii) Fine-tuning methods perform well with just one image, as shown in Appendix A.
C.2 Which technical route performs better?
Our comprehensive analysis reveals distinct advantages of different technical routes across various generation scenarios. For concept preservation, encoder-based methods demonstrate superior performance in human-centric generation, particularly in preserving precise facial features where fine-tuning approaches often struggle. This aligns with the recent trend in Identity-Preserving Generation tasks. However, fine-tuning methods show stronger capabilities in handling common objects and standardized appearances, while encoder-based approaches provide more consistent performance across diverse object types.
The choice of base model significantly impacts generation quality. SDXL exhibits better concept preservation and prompt following capabilities compared to SD v1.5, primarily due to its more sophisticated text encoder architecture and diverse training data. Regarding prompt adherence, fine-tuning methods generally outperform encoder-based approaches by better preserving the original text-image conditional distribution through direct learning from image-text pairs. In contrast, encoder-based methods may face challenges in maintaining precise prompt alignment due to potential disruptions during feature injection.
These findings suggest that the optimal technical route depends heavily on the specific application scenario. For applications prioritizing human identity preservation, encoder-based methods are preferable. However, for use cases requiring precise prompt following or handling common objects, fine-tuning approaches might be more suitable. This understanding helps guide future research directions in personalized image generation.
Appendix D Limitation & Future Work
Human-aligned evaluation & benchmarking is an emerging but challenging direction, and we have only made preliminary attempts at personalized image generation. Moreover, our evaluation results heavily rely on the advancements of multimodal large language models and require carefully designed system prompts. We believe that as visual world models continue to develop, the evaluation performance will be further optimized. Our future work will focus on more applications with human-aligned evaluation, such as multimodal intelligence (Wei et al., 2024a; Zhu et al., 2025; Wei et al., 2024d; c; Zhu et al., 2024; Liu et al., 2024; Chen et al., 2024; Wei et al., 2024b; Gao et al., 2024a; b), audio generation (Gao et al., 2024c), 3D generation (Poole et al., 2023; Qi et al., 2023; Liu et al., 2023c; Gafni et al., 2022; Gao et al., 2023), video generation (Ho et al., 2022; Singer et al., 2023; Blattmann et al., 2023; Zhou et al., 2025), autonomous driving (Han et al., 2024; Li et al., 2023c; Zhang et al., 2023b), and even embodied visual intelligence (Goyal et al., 2022; Liang et al., 2023a; Brohan et al., 2023; Driess et al., 2023; Qi et al., 2025; Zhao et al., 2024; He et al., 2024b; Yu et al., 2024; 2025).